# SINDI: an Efficient Index for Approximate Maximum Inner Product Search on Sparse Vectors

### v1

@@ 5.5. Scalability
To further evaluate the scalability of the Sindi algorithm, we conducted a multi-threaded performance test on two large-scale datasets: SPLADE-FULL and AntSparse-10M. We measured QPS at different recall targets ($\mathrm{Recall}@50\in\{0.91,0.95,0.999\}$) while varying the number of CPU cores from 2 to 10, as shown in Figure 14. On AntSparse-10M at $\mathrm{Recall}@50=0.90$, using 2 CPU cores yields 1979.49 QPS (approximately 989.75 QPS per core), while using 10 cores achieves 8374.01 QPS (approximately 837.40 QPS per core). The per-core efficiency remains high, dropping by less than 16% when scaling from 2 to 10 cores. Similar scaling behavior is observed for SPLADE-FULL, confirming that Sindi effectively utilizes available CPU cores with minimal parallelization overhead.
These results demonstrate that Sindi maintains high multi-core efficiency across datasets and accuracy levels, making it suitable for deployment in scenarios requiring both high recall and high throughput.
@@ 6. Related Work
() SPLADE-FULL
(a) AntSparse-10M
Existing methods for MIPS on sparse vectors adopt diverse index structures, including inverted index-, graph-, and hash-based designs. Inverted index-based methods, such as Seismic (seismic, ), build postings for each dimension and prune low-value entries to reduce the candidate set. They cluster postings into blocks to skip irrelevant candidates. However, as each vector appears in multiple postings, value retrieval still requires random access to scattered data.
Graph-based methods, such as Hnsw-based PyAnn(hnsw, ), organize vectors as nodes in a proximity graph. At query time, greedy graph traversal identifies neighbors, but vector sparsity leads to weak connectivity and frequent random memory accesses.
Hash-based methods, like Sosia (sosia, ), transform sparse vectors into sets (via SOS) and use min-hash to estimate Jaccard similarity. Multiple hash functions improve accuracy but scatter vector storage across buckets, limiting locality and making direct computation in postings impractical.
Bmp (bmp, ; sp, ) also stores non-zero values in postings and prunes with block-level maximum value vectors. While this avoids ID lookups, its fine-grained partitioning leads to excessive block evaluations, degrading efficiency. Our method retains value-stored postings without requiring costly block-level filtering.
@@ 7. Conclusion
In this work, we propose Sindi, an inverted index for sparse vectors that eliminates redundant distance computations. By storing term weights directly in the postings, Sindi removes both the ID lookup and random memory access overhead in distance computation, and leverages SIMD to fully exploit CPU parallelism. It further introduces Mass Ratio Pruning that preserves maximum inner-product accuracy while enabling scalable approximate search on large datasets. Experiments on multilingual, multi-scale real-world datasets demonstrate that Sindi achieves state-of-the-art performance.

### v2

@@ 5.5. Scalability
To further evaluate the scalability of the Sindi algorithm, we conducted a multi-threaded performance test on two large-scale datasets: SPLADE-FULL and AntSparse-10M. We measured QPS at different recall targets, with $\mathrm{Recall}@50\in\{0.95,0.97,0.99\}$ for SPLADE-FULL and $\mathrm{Recall}@50\in\{0.90,0.95,0.99\}$ for AntSparse-10M, while varying the number of CPU cores from 2 to 10, as shown in Figure 14.
On AntSparse-10M at $\mathrm{Recall}@50=0.90$, using 2 CPU cores yields 1979.49 QPS ($\approx 989.75$ QPS per core), whereas 10 cores achieve 8374.01 QPS ($\approx 837.40$ QPS per core), indicating per-core efficiency drops by less than 16% when scaling from 2 to 10 cores. On SPLADE-FULL at $\mathrm{Recall}@50=0.99$, using 2 CPU cores yields 453.46 QPS ($\approx 226.73$ QPS per core), whereas 10 cores achieve 2142.05 QPS ($\approx 214.21$ QPS per core), corresponding to a per-core efficiency drop of about 5.5%. Similar scaling behavior is observed across other recall targets for both datasets.
These results show that Sindi maintains high multi-core efficiency across datasets and accuracy targets, confirming its suitability for deployment in scenarios requiring both high recall and high throughput, with minimal parallelization overhead.
@@ 6. Related Work
Existing methods for MIPS on sparse vectors employ various index structures, including inverted index, graph, and hash-based designs. Inverted index-based methods, such as Seismic (Bruch et al., 2024), maintain postings for each dimension and prune low-value entries, often clustering postings into blocks, but each vector may appear in multiple postings, causing random access to scattered memory and reducing locality.
Graph-based methods, such as the HNSW-based PyANNS (Malkov and Yashunin, 2020), organize vectors as nodes in a proximity graph, but vector sparsity weakens connectivity and greedy traversal still incurs frequent random memory accesses.
Hash-based methods, like Sosia (Zhao et al., 2024), transform sparse vectors into sets via the SOS representation and apply min-hash to estimate Jaccard similarity. Multiple hash functions improve accuracy but scatter storage across many buckets, hurting locality and making direct computation inefficient.
BMP (Mallia et al., 2024; Carlson et al., 2025) stores non-zero values in postings and prunes them using block-level maximum value vectors. While this avoids common-dimension lookups, overly fine-grained block partitioning increases the number of candidate blocks to evaluate, lengthening query time. In contrast, our method keeps value-stored postings without incurring such overhead.
@@ 7. Conclusion
In this work, we propose Sindi, an inverted index for sparse vectors that eliminates redundant inner-product computations. By storing non-zero entries directly in the postings, Sindi removes both document ID lookups and random memory accesses, and leverages SIMD acceleration to maximize CPU parallelism. It further introduces Mass Ratio Pruning, which effectively preserves high-value entries, and a reordering strategy whose refinement step ensures high accuracy. Experiments on multilingual, multi-scale real-world datasets demonstrate that Sindi delivers state-of-the-art performance in both recall and throughput.